# Imaging the Photochemistry of Cyclobutanone using Ultrafast Electron Diffraction: Experimental Results

### v1

@@ Abstract
We investigated the ultrafast structural dynamics of cyclobutanone following photoexcitation at nm using gas-phase megaelectronvolt ultrafast electron diffraction. Our investigation complements the simulation studies of the same process within this special issue. It provides information about both electronic state population and structural dynamics through well-separable inelastic and elastic electron scattering signatures. We observe the depopulation of the photoexcited S2 state of cyclobutanone with n3s Rydberg character through its inelastic electron scattering signature with a time constant of ps towards the S1 state. The S1 state population undergoes ring-opening via a Norrish Type-I reaction, likely while passing through a conical intersection with S0. The corresponding structural changes can be tracked by elastic electron scattering signatures. These changes appear with a delay of ps with respect the initial photoexcitation, which is less than the S2 depopulation time constant. This behavior provides evidence for the ballistic nature of the ring-opening once the S1 state is reached. The resulting biradical species react further within ps via two rival fragmentation channels yielding ketene and ethylene, or propene and carbon monoxide. Our study showcases both the value of gas-phase ultrafast diffraction studies as an experimental benchmark for nonadiabatic dynamics simulation methods and the limits in the interpretation of such experimental data without comparison to such simulations.
@@ I Introduction
The ultrafast dynamics of CB after excitation at nm have been investigated by a combination of time-resolved photoelectron spectroscopy and time-resolved mass spectrometry.Kuhlman, Sølling, and Møller (2012) The time-dependences of the parent ion and a fragment mass corresponding to the ketene and the cyclopropane or propene fragments were fitted to biexponential decays with time constants of 0.08 ps and 0.74 ps for the parent, and 0.12 ps and 0.79 ps for the fragment, respectively. The time-resolved photoelectron spectroscopy study identified a signature displaying a biexponential decay with time constants of 0.31 ps and 0.74 ps, suggesting depopulation of the n3s state on an ultrafast timescale via internal conversion to the n state. The decay of the photoelectron signature was observed to be modulated both in position and intensity by an oscillatory component with a period of 0.94 ps.
The evidence for ultrafast nonadiabatic dynamics between states of different electronic character and the rich photochemistry, with several intermediates and products, make CB an ideal benchmark molecule for a challenge for theoreticians to predict non-adiabatic photochemical dynamics in advance of an experiment, as pursued in this special issue.Bennett et al. (2024); Eng, Rankine, and Penfold (2024); Hait et al. (2024); Hutton et al. (2024); Jaiswal et al. (2024); Janoš et al. (2024); Lawrence et al. (2024); Makhov et al. (2024); Martín Santa Daría et al. (2024); Miao, Diemer, and Mitrić (2024); Suchan et al. (2024); Miller et al. (2024); Mukherjee et al. (2024); Peng, Liu, and Lan (2024); Vindel-Zandbergen and González-Vázquez (2024) In the following, we will discuss the results of our UED study of the photochemical dynamics of CB after excitation to the n3s state at nm. Since our study is the experimental complement to the companion simulation studies, we seek to isolate the analysis of the experimental data as much as possible from the simulation results. Therefore, we will use only a basic level of quantum chemical calculations (geometry optimizations in the ground state at the density functional theory (DFT) level) resulting in an interpretation of the experimental results, which is independent from more extensive simulations.
@@ II Methods
Cyclobutanone (purity 99 %) is purchased from Sigma Millipore and used without further purification. We employ a static-filled 3 mm flow cell (500 m orifices, sample at room temperature) in combination with a repetition rate of 360 Hz. Diffracted electrons are detected by a combination of a phosphor screen and an electron multiplying charge-coupled device (EMCCD) camera. Based on the relative levels of static and dynamic signal, we estimate that we excite about 0.3 % of the molecules (see the Supplementary Information (SI), sect. S1 for details). Therefore, we do not expect to observe any contributions from multiphoton absorption. Time-dependent diffraction is measured at a series of time delay points between -4.7 ps and 2.1 ps in each scan, where negative delays correspond to the electron probe arriving before the optical pump in the sample. The separation between time delay points is 0.05 ps, except for the earliest and latest delay points where it is considerably larger. At each time delay point, we integrate the diffraction signal for 10 s (3600 shots). The full data set includes 393 such scans of all delays. The processing of the raw diffraction patterns obtained is described in detail in the SI, sect. S2.
@@ IV.2 The origin of the feature and the onset of the structural dynamics
Without more advanced simulation of the observable, we cannot exclude the possibility that the feature also contains contributions from population in the n state at later delays. However, the fitted decay time constant (0.29 ps, more details in SI sect. S5) of the feature agrees very well with the faster time constant of the biexponential fit to the n3s feature in the time-resolved photoelectron spectra from Ref. Kuhlman, Sølling, and Møller (2012). Therefore, the feature is likely predominantly caused by population in the n3s state.
@@ IV.3 Delay and shift in the feature
The onset of the feature does not give evidence for the specific nature of the structural change, since it is shared by both the primary and secondary products. The slower onset of the feature and its shift to lower -values provide more conclusive insights into this issue. The signature of the biradical geometry shows a positive contribution to higher -values with respect to the secondary products. A transformation from the primary to the secondary products would result in the observed shifts in both the and features to lower -values.
The assignment of the experimental signature in Fig. 4 c) to the biradical product could be more obvious after real-space transformation. However, the feature dominates the signal at these early delays. Due to its suspected inelastic scattering nature, a real-space transformation would not yield meaningful results.
One would intuitively expect to observe the onset of the feature to be co-timed with the onset of the feature at slightly higher -values. Instead, it slowly increases in intensity while also shifting in . This seeming disagreement is resolved by considering the much more fluxual nature (see sect. IV.1) of the primary photoproduct compared to the secondary products. In the comparison of the signatures of different possible geometries (see the SI, sect. S3), the peak positions in the and regions show the largest variance, while the peak position in the region is much more stable. Based on this observation, we expect that the experimental signature of the primary photoproduct is strongly blurred and, therefore, rather weak in the region, which could explain the slow onset of the feature. Additionally, due to its transitory nature, its population at all but the early pump-probe delays can be expected to be low contributing to the observed weakness in the region.
@@ IV.4 Proposed reaction mechanism based on the UED results
The lifetime of the feature (0.29 0.2 ps) is larger than the difference in onset between the and the features. This finding can be explained by the depopulation of the n3s state being preceded by an induction period of () ps (the delay between and onset). During the induction period, the n3s population evolves from the Franck-Condon region to the conical intersection with the n state. This evolution does not represent a sufficiently large enough structural change to result in a measurable signal. Having traversed the n3s/n conical intersection, molecules on the n potential undergo quasi-instantaneous C1-C4 bond extension and access the conical intersection of the n state with the ground state. Internal conversion through this conical intrsection reults in vibrationally hot ground state molecules.

### v2

@@ Abstract
We investigated the ultrafast structural dynamics of cyclobutanone following photoexcitation at nm using gas-phase megaelectronvolt ultrafast electron diffraction. Our investigation complements the simulation studies of the same process within this special issue. It provides information about both electronic state population and structural dynamics through well-separable inelastic and elastic electron scattering signatures. We observe the depopulation of the photoexcited S2 state of cyclobutanone with n3s Rydberg character through its inelastic electron scattering signature with a time constant of ps towards the S1 state. The S1 state population undergoes ring-opening via a Norrish Type-I reaction, likely while passing through a conical intersection with S0. The corresponding structural changes can be tracked by elastic electron scattering signatures. These changes appear with a delay of ps with respect to the initial photoexcitation, which is less than the S2 depopulation time constant. This behavior provides evidence for the ballistic nature of the ring-opening once the S1 state is reached. The resulting biradical species react further within ps via two rival fragmentation channels yielding ketene and ethylene, or propene and carbon monoxide. Our study showcases both the value of gas-phase ultrafast diffraction studies as an experimental benchmark for nonadiabatic dynamics simulation methods and the limits in the interpretation of such experimental data without comparison to such simulations.
@@ I Introduction
The ultrafast dynamics of CB after excitation at nm have been investigated by a combination of time-resolved photoelectron spectroscopy and time-resolved mass spectrometry.Kuhlman, Sølling, and Møller (2012) The time-dependences of the parent ion and a fragment mass corresponding to the ketene and the cyclopropane or propene fragments were fitted to biexponential decays with time constants of 0.08 ps and 0.74 ps for the parent, and 0.12 ps and 0.79 ps for the fragment, respectively. The time-resolved photoelectron spectroscopy study identified a signature displaying a biexponential decay with time constants of 0.31 ps and 0.74 ps, suggesting depopulation of the n3s state on an ultrafast timescale via internal conversion to the n state. The decay of the photoelectron signature was observed to be modulated both in position and intensity by an oscillatory component with a period of 0.94 ps, which could be associated with motion along a ring-puckering mode in the n3s state.Kuhlman et al. (2013)
The evidence for ultrafast nonadiabatic dynamics between states of different electronic character and the rich photochemistry, with several intermediates and products, make CB an ideal benchmark molecule for a challenge for theoreticians to predict non-adiabatic photochemical dynamics in advance of an experiment, as pursued in this special issue.Bennett et al. (2024); Eng, Rankine, and Penfold (2024); Hait et al. (2024); Hutton et al. (2024); Jaiswal et al. (2024); Janoš et al. (2024); Lawrence et al. (2024); Makhov et al. (2024); Martín Santa Daría et al. (2024); Miao, Diemer, and Mitrić (2024); Suchan et al. (2024); Miller et al. (2024); Mukherjee et al. (2024); Peng, Liu, and Lan (2024); Vindel-Zandbergen and González-Vázquez (2024)
Thus, the purpose of this study is to provide (1) an experimental benchmark to the predictions published in this special issue and (2) information about the mechanism and timescale of the photochemical reaction dynamics of CB following photoexcitation. In the following, we will discuss the results of our UED study of the photochemical dynamics of CB after excitation to the n3s state at nm. Since our study is the experimental complement to the companion simulation studies, we seek to isolate the analysis of the experimental data as much as possible from the simulation results. Therefore, we will use only a basic level of quantum chemical calculations (geometry optimizations in the ground state at the density functional theory (DFT) level) resulting in an interpretation of the experimental results, which is independent from more extensive simulations.
@@ II Methods
Cyclobutanone (purity 99 %) is purchased from Sigma Millipore and used without further purification. The compound has a vapor pressure of 43 torr at room temperature.Benson and Kistiakowsky (1942) This pressure is reduced with the help of a flow controller (MKS) to 1 torr. The resulting low-pressure gas is delivered into a static-filler 3 mm flow cell with 500 m diameter entrance and exit holes for the pump and probe beams.Lin et al. (2023) The experiment is performed at a repetition rate of 360 Hz. Diffracted electrons are detected by a combination of a phosphor screen and an electron multiplying charge-coupled device (EMCCD) camera. Based on the relative levels of static and dynamic signal, we estimate that we excite about 0.3 % of the molecules (see the Supplementary Information (SI), sect. S1 for details). Therefore, we do not expect to observe any contributions from multiphoton absorption. Time-dependent diffraction is measured at a series of time delay points between -4.7 ps and 2.1 ps in each scan, where negative delays correspond to the electron probe arriving before the optical pump in the sample. The separation between time delay points is 0.05 ps, except for the earliest and latest delay points where it is considerably larger. At each time delay point, we integrate the diffraction signal for 10 s (3600 shots). The full data set includes 393 such scans of all delays. The processing of the raw diffraction patterns obtained is described in detail in the SI, sect. S2.
@@ IV.2 The origin of the feature and the onset of the structural dynamics
Without more advanced simulation of the observable, we cannot exclude the possibility that the feature also contains contributions from population in the n state at later delays. However, the fitted decay time constant (0.29 ps, more details in SI sect. S5) of the feature agrees very well with the faster of the two time constants of the biexponential fit (0.31 ps and 0.74 ps) to the n3s feature in the time-resolved photoelectron spectra from Ref. Kuhlman, Sølling, and Møller (2012). Therefore, the feature is likely predominantly caused by population in the n3s state.
@@ IV.3 Delay and shift in the feature
The onset of the feature does not give evidence for the specific nature of the structural change, since it is shared by both the primary and secondary products. The slower onset of the feature and its shift to lower -values provide more conclusive insights into this issue. The signature of the biradical geometry shows a positive contribution to higher -values with respect to the secondary products in the beta region. Thus, a transformation from the biradical to the secondary products would result in the observed shift in the feature to lower -values. Hence, the experimental signature at delays between 0.2 ps and 0.5 ps in Fig. 4 c) can be assigned to the biradical photoproduct. This assignment could be more obvious after real-space transformation. However, the feature dominates the signal at these early delays. Due to its suspected inelastic scattering nature, a real-space transformation would not yield meaningful results.
One would intuitively expect to observe the onset of the feature to be co-timed with the onset of the feature at slightly higher -values. Instead, it slowly increases in intensity while also shifting in . This behavior can be explained by considering the much less well-defined geometry (see sect. IV.1) of the primary photoproduct compared to the secondary products. In the comparison of the signatures of different possible biradical geometries (see the SI, sect. S3), the peak positions in the and regions show the largest variance, while the peak position in the region is much more stable. Based on this observation, we expect that the experimental signature of the primary photoproduct is much less pronounced than the secondary products in the region, which could explain the slow onset of the feature. Additionally, due to its transitory nature, its population at all but the early pump-probe delays can be expected to be low contributing to the observed weakness in the region.
@@ IV.4 Proposed reaction mechanism based on the UED results
The lifetime of the feature (0.29 0.2 ps) is larger than the difference in onset between the and the features. However, according to previous time-resolved photoelectron spectroscopy results, the main structural motion enabling depopulation of the n3s state is ring-puckering,Kuhlman et al. (2013) which can be expected to provide a negligibly weak signature in the difference diffraction patterns. Thus, having traversed the n3s/n conical intersection, molecules on the n potential must undergo quasi-instantaneous (C1-C4) bond extension and access the conical intersection of the n state with the ground state. Accordingly, the diffraction patterns show the onset of structural dynamics in the n state (onset of the feature) while the population dynamics out of the n3s state (decay of the feature) are still ongoing. Internal conversion through the conical intersection with the ground state results in vibrationally hot ground state molecules.